OpenStudy (anonymous):

Simplify expression. Use positive exponents. a\[a^4b-^3/ab-^2\] Can someone help me solve this and tell me the steps on how to do so? I don't get this at all

OpenStudy (anonymous):

\[a^4b-^3/ab-^2\]

OpenStudy (anonymous):

well whenever there is a negative exponent, like b^-3, it would then become 1/b^3

OpenStudy (anonymous):

\[1/b^3\]?

OpenStudy (danjs):

\[\frac{ a^4*b ^{-3} }{ a*b ^{-2} } = \frac{ a ^{4-1} }{ b ^{-2+3} }\]

OpenStudy (danjs):

When you change a base to a power to the other side of the fraction, you change the sign on the power

OpenStudy (danjs):

and multiplying like bases to powers, you add the powers

OpenStudy (anonymous):

@DanJS thank you so much :~}

OpenStudy (danjs):

IE a^4/a^1 = a^(4-1) = a^3

OpenStudy (danjs):

no prob, just have to remember the 4 or 5 power rules

OpenStudy (anonymous):

@DanJS what type of problem is this? like is it rational or just and equatiom?

OpenStudy (anonymous):

\[\frac{ 1a^4 }{ b ^{3} } / \frac{ 1 }{ ab ^2 }\] so then you would basically multiply by the recipricol, ending up in \[\frac{ 1a^4 }{ b ^{3} } * \frac{ ab^2 }{ 1}\] you would then end up having \[\frac{ a ^{5} b ^{2} }{ b^3} \] and according to the power rules, you would end up canceling out \[b ^{2}\], ending up with \[ \frac{ a ^{5} }{ b }\]

OpenStudy (anonymous):

this is really just an expression problem so you aren't really solving for an actual numeric answer

OpenStudy (danjs):

\[\frac{ a^4*b ^{-3} }{ a*b ^{-2} } = \frac{ a^3 }{ b }\]

OpenStudy (danjs):

from the above

OpenStudy (anonymous):

if you wanna check my work DanJS please do:)

OpenStudy (anonymous):

WAIT don't look at my answer, I just realized I read your question wrong! Go with DanJS's answer:))